# Robust Market Interventions

### v1

@@ 7.4. Nonlinear demand
When demand is not locally linear, the pass-through of cost shocks depends not only on the Slutsky matrix (which is the Jacobian of demand) but also on the Hessian of the demand function, the matrix whose $(i,j)$ entry is $\partial^{2}q_{i}(\bm{p})/\partial p_{i}\partial p_{j}$ (see, e.g., Miklos-Thal and Shaffer (2021)). Thus, our calculations would change, and there is no guarantee that our linear tax/subsidy interventions based only on the Slutsky matrix at the status quo, would perform as they do in the linear model.
However, our main result can be extended once we allow the authority to use nonlinear interventions, i.e., to commit a vector of functions specifying a payment to each producer $i$ as a function of all prices and quantities realized after the intervention. With this broader set of instruments, the authority can use nonlinear rebates based on post-intervention quantities to reduce the problem to the one we have studied. The key idea is to effectively linearize the demand the firms face around the status quo—i.e., use transfers to make up the difference between realized demand and a linear demand function. Once demand has been “linearized” in this way, the problem that firms face becomes equivalent to the one we have studied and we can use the results developed to design per-unit tax/subsidy interventions with desirable welfare properties. If we assume that the curvature of the demand of each product is locally bounded by a known constant, the payments needed to linearize demand can be bounded by a small fraction of the first-order gains of an intervention, so our welfare guarantees go through. Such assumptions on curvature also allow us specify concrete sizes of interventions that achieve a given level of welfare gain, rather than just characterizing the behavior of derivatives.
@@ A.2. Proof of 1
Our first main lemma, which we will prove shortly, will assert that $\dot{S}$ converges in probability to $1$. The challenge in proving such a result is that the actual expenditure depends on true quantities, whereas the intervention is built based on estimated quantities projected onto an “estimated” eigenspace of $\widehat{\bm{D}}$. We begin with a technical result that will be key to controlling the differences between actual and estimated objects. It relies on the Davis–Kahan Theorem, which is important for our analysis.
@@
The Davis–Kahan Theorem guarantees that, for any $\eta>0$, there exists $N_{1}(\eta)$ such that for all $n>N_{1}(\eta)$, with probability greater than or equal to $1-\eta$, two key properties hold. First, every eigenvalue in the set $\Lambda(n)$, which is nonempty by the recoverable structure assumption, has a corresponding eigenvalue within distance $O(b(n))$, and therefore in $\widehat{\Lambda}(n)$. Second,
$\|P_{\underline{L}(n)}^{\perp}P_{\widehat{L}(n)}\|\leq\frac{2\|\bm{E}\|}{\text{gap}}$ (14)
where gap is the minimum distance between some eigenvalue of $\widehat{\bm{D}}$ in $\widehat{\Lambda}(n)$ and some eigenvalue of $\bm{D}$ not contained in $\underline{\Lambda}(n)$. This gap is at least $\widehat{M}(n)-\underline{M}(n)\gg b(n)$. Thus, increasing $N_{1}(\eta)$ if necessary, we conclude the statement of the lemma. ∎
@@
Now, to prove the theorem, we use 2 to write:
@@
This follows from 3 and the fact that by construction, $\bm{\sigma}\in\widehat{L}(n)$. ∎
@@
$|\dot{W}-\dot{S}|={\sum_{\lambda_{\ell}\in\underline{\Lambda}(n)}\left[\frac{|\lambda_{\ell}|}{1+|\lambda_{\ell}|}-1\right](\bm{u}^{\ell}\cdot\bm{q}^{0})(\bm{u}^{\ell}\cdot\bm{\sigma})}$
@@ A.3. Proof of 2
Part 1: We construct an environment (i.e, a set of market states $\bm{\Theta}$ and, for each $\bm{\theta}\in\bm{\Theta}$ a distribution $\mu_{\bm{\theta}}$ of $(\widehat{\bm{D}},\hat{\bm{q}}^{0})$) where no intervention robustly increases total surplus, but where an omniscient authority could achieve total surplus gains $\dot{P}+\dot{C}=2\dot{S}$ with $\dot{C}\geq 0$.
@@
In this proof it will be useful to define random variables on the extended reals {$\mathbb{R}\cup\{-\infty,\infty\}$, which allows us to take $\varepsilon_{i}=\infty$ for all $i$ [^30] In 4, we may take $\overline{V}=\infty$; this is immaterial since that assumption plays a role only in our positive results. Since $\bm{D}$ is deterministic, we can take $\bm{E}$ to be the zero matrix. Note that $\bm{D}$ is known and the signal about $\bm{q}^{0}$ is always $\infty$ in every entry, so that any intervention rule implements a single intervention—call it $\bm{\sigma}^{*}$.
@@
For any intervention $\bm{\sigma}$, 2 states that the change in total surplus is given by:
@@
On the other hand, by 3 in Appendix B, any surplus outcome $(\dot{P},\dot{C},\dot{S})$ satisfying $\frac{1}{2}\dot{P}+\dot{C}=\dot{S}$ can be achieved.
Part 2: Note the setting here is different from Part 1: in that case we were constructing an environment without aggregate structure, while here we will give an example of an environment with aggregate structure where no intervention can robustly increase consumer surplus.
@@
We will study the distribution of this random variable for any fixed $\bm{\sigma}$, under the prior. We claim that $\dot{C}_{\geq 2}$ has the same distribution as $-\dot{C}_{\geq 2}$. This is because in the above expression only $\bm{q}^{0}$ is random, and $\bm{u}^{\ell}\cdot\bm{q}^{0}$ are independently equal to plus or minus the same constant by definition. Thus, for any $\epsilon>0$,
@@
On the other hand, again by 3 in Appendix B, any surplus outcome $(\dot{P},\dot{C},\dot{S})$ satisfying $\frac{1}{2}\dot{P}+\dot{C}=\dot{S}$ can be achieved.
@@
For simplicity this proof has worked with a case where all uncertainty comes from $\bm{q}^{0}$, and the projections of $\bm{q}^{0}$ onto various low eigenvectors are small. A stronger negative result can be obtained by elaborating the construction a bit—making the $\lambda_{\ell}$ different from one another, so that the eigenspaces orthogonal to $\bm{1}$ are not deterministic. Then, by showing that the eigenvectors are often impossible to recover with any precision, we can obtain a similar conclusion even if all the projections $\bm{u}^{\ell}\cdot\bm{q}^{0}$ are bounded below.
@@ Appendix B Complete information
First, consider the case in which products are independent. In this case, the effect of a subsidy to firm $i$ and a subsidy to firm $j$ to surpluses are separable. So it is sufficient to prove the result for an intervention which only susidizes one firm: take $\bm{\sigma}$ with $\sigma_{j}=0$ for all $j\neq 1$ and $\sigma_{1}>0$, so that $\dot{S}=\sigma_{1}q_{1}^{0}$. Note the basis diagonalizing $\bm{D}$ is the standard basis, and using the formulas in 2 with $\lambda_{\ell}=-1$ for all $\ell$, we conclude $\dot{C}=\sigma q_{1}^{0}/2$ and $\dot{P}=\sigma q_{1}^{0}$, as claimed.
@@ Appendix C A noise structure under which 4 holds
For simplicity we assume that all households share a single representative utility function for goods and that the number of households exceeds $n^{2}$, where $n$ is the number of firms. Moreover, we assume that, for all $i$ and $j$, the entries $\underline{D}_{ij}$ are uniformly bounded, i.e., $|\underline{D}_{ij}|\leq\underline{D}_{\operatornamewithlimits{max}}<\infty$ for some positive constant $\underline{D}_{\operatornamewithlimits{max}}$. Also assume the true diagonal entries $\underline{D}_{ii}$ are bounded away from 00: $0<\underline{D}_{\operatornamewithlimits{min}}\leq|\underline{D}_{ii}|$ for all $i$. The constants $\underline{D}_{\operatornamewithlimits{min}}$ and $\underline{D}_{\operatornamewithlimits{max}}$ do not depend on $i,j$, or $n$.
@@
We assume that the $F_{ij}$ are mean-zero and independent across $(i,j)$, that their support is uniformly bounded, and that the matrix $\bm{F}$ is symmetric. [^32] Since $\bm{D}$ is a symmetric matrix, we view the paramters as being the numbers on or above the diagonal. Let $\overline{V}$ be an upper bound on $\text{Var}[F_{ij}]$ that does not depend on $i,j$, or $n$.
@@
By standard results on the expected spectral norm of symmetric random matrices with independent entries (bai1988necessary), we have that $\|\bm{H}\|]=O_{p}(\sqrt{n})$. Under our assumption that $\bm{F}$ has uniformly bounded support, this can be extended (dallaporta2012eigenvalue, Theorem 2) to the statement that $\mathbb{E}[\|\bm{H}\|]=O(\sqrt{n})$.

### v2

@@ 7.4. Nonlinear demand
When demand is not locally linear, the pass-through of marginal cost shocks depends not only on the Slutsky matrix (which is the Jacobian of demand) but also on the Hessian of the demand function, the matrix whose $(i,j)$ entry is $\partial^{2}q_{i}(\bm{p})/\partial p_{i}\partial p_{j}$ (see, e.g., Miklos-Thal and Shaffer (2021)). Thus, our calculations would change, and there is no guarantee that our linear tax/subsidy interventions (based only on the Slutsky matrix at the status quo) would perform as they do in the linear model.
However, our main result can be extended once we allow the authority to use nonlinear interventions, i.e., to commit a vector of functions specifying a payment to each producer $i$ as a function of all prices and quantities realized after the intervention. With this broader set of instruments, the authority can use nonlinear rebates based on post-intervention quantities to reduce the problem to the one we have studied. The key idea is to effectively linearize the demand the firms face around the status quo by using transfers to make up the difference between realized demand and a linear demand function. Once demand has been “linearized” in this way, the problem that firms face becomes equivalent to the one we have studied and we can use the results developed to design per-unit tax/subsidy interventions with desirable welfare properties. If we assume that the curvature of the demand of each product is locally bounded by a known constant, the payments needed to linearize demand can be bounded by a small fraction of the first-order gains of an intervention, so our welfare guarantees remain valid. Such assumptions on curvature also allow us to specify concrete sizes of interventions that achieve a given level of welfare gain, rather than just characterizing the behavior of derivatives.
@@ A.2. Proof of 1
Fix an $n$ and a set of types $\{1,2,\ldots,m\}$, with the number of products of type $t$ being $N(t)$.
The matrix $\bm{D}$ can be written as
$\bm{D}=\bm{A}+(1+\gamma(n))\bm{I}$
where $\bm{A}$ is a block matrix, in the sense that $A_{ij}$ depends only on $k(i)$ and $k(j)$.
The nonzero eigenvalues of $\bm{A}$ are the same as those of the $m$-by-$m$ matrix $\widetilde{\bm{A}}$ whose $(t,t^{\prime})$ entry is
$\widetilde{A}_{tt^{\prime}}=\gamma(n)\sqrt{N(t)N(t^{\prime})}\bar{D}_{tt^{\prime}}.$
There is some $t$ so that $N(t)\geq n/|M|$. Thus,
$|\widetilde{A}_{tt^{\prime}}|\geq\frac{\gamma(n)}{|M|}n|\bar{D}_{tt}|=\frac{\gamma(n)}{|M|}n,$
recalling $\bar{\bm{D}}$ has diagonal $-1$. By the Courant–Fisher Theorem, this is a lower bound on the magnitude of the largest eigenvalue $\widetilde{\bm{A}}$. Since the adjustment coming from the scaled identity matrix is of lower order, $\bm{D}$ has an eigenvalue at least $\gamma(n)n/2|M|$ in magnitude.
Now let $\widetilde{\bm{u}}^{1}$ be the corresponding eigenvector of $\bm{\widetilde{A}}$ normalized to lie in the unit ball, and $\widetilde{\bm{v}}^{1}$ be its limit as $n\to\infty$. This limit depends only on the limit $\widetilde{\bm{A}}/n$, which exists by our assumption that $N(t)/n$ converges for each $n$. The limit projection of $\bm{q}^{0}$ onto $\bm{u}^{1}$ depends only on $\bar{\bm{q}}$ and $\bm{v}^{1}$, and for generic $\bm{q}$ this is nonzero.
A.3. Proof of 1
@@
Our first main lemma, which we will prove shortly, will assert that $\dot{S}$ converges in probability to $1$. The challenge in proving such a result is that the actual expenditure depends on true quantities, whereas the intervention is built based on estimated quantities projected onto an “estimated” eigenspace of $\widehat{\bm{D}}$. We begin with a technical result that will be key to controlling the differences between actual and estimated objects. It relies on the Davis–Kahan Theorem.
@@
The Davis–Kahan Theorem guarantees that, for any $\eta>0$, there exists $N_{1}(\eta)$ such that for all $n>N_{1}(\eta)$, with probability greater than or equal to $1-\eta$, two properties hold. First, every eigenvalue in the set $\Lambda(n)$, which is nonempty by the recoverable structure assumption, has a corresponding eigenvalue within distance $O(b(n))$, and therefore in $\widehat{\Lambda}(n)$. Second,
$\|P_{\underline{L}(n)}^{\perp}P_{\widehat{L}(n)}\|\leq\frac{2\|\bm{E}\|}{g},$ (17)
where $g$, the “gap,” is the minimum distance between some eigenvalue of $\widehat{\bm{D}}$ in $\widehat{\Lambda}(n)$ and some eigenvalue of $\bm{D}$ not contained in $\underline{\Lambda}(n)$. This gap is at least $\widehat{M}(n)-\underline{M}(n)\gg b(n)$. Thus, increasing $N_{1}(\eta)$ if necessary, we conclude the statement of the lemma. ∎
@@
Now, to prove the theorem, we use 3 to write:
@@
This follows from 4 and the fact that by construction, $\bm{\sigma}\in\widehat{L}(n)$. ∎
@@
$|\dot{W}-\dot{S}|={\sum_{\lambda_{\ell}\in\underline{\Lambda}(n)}\left[\frac{|\lambda_{\ell}|}{1+|\lambda_{\ell}|}-1\right](\bm{u}^{\ell}\cdot\bm{q}^{0})(\bm{u}^{\ell}\cdot\bm{\sigma})}.$
@@ A.4. Proof of 3
Part 1: We construct an environment (i.e., a set of market states $\bm{\Theta}$ and, for each $\bm{\theta}\in\bm{\Theta}$ a distribution $\mu_{\bm{\theta}}$ of $(\widehat{\bm{D}},\hat{\bm{q}}^{0})$) where no intervention robustly increases total surplus, but where an omniscient authority could achieve total surplus gains $\dot{P}+\dot{C}=2\dot{S}$ with $\dot{C}\geq 0$.
@@
In this proof, it will be useful to define random variables on the extended reals {$\mathbb{R}\cup\{-\infty,\infty\}$, which allows us to take $\varepsilon_{i}=\infty$ for all $i$ [^36] In 4, we may take $\overline{V}=\infty$; this is immaterial since that assumption plays a role only in our positive results. Since $\bm{D}$ is deterministic, we can take $\bm{E}$ to be the zero matrix. Note that $\bm{D}$ is known and the signal about $\bm{q}^{0}$ is always $\infty$ in every entry, so that any intervention rule implements a single intervention—call it $\bm{\sigma}^{*}$.
@@
For any intervention $\bm{\sigma}$, 3 states that the change in total surplus is given by:
@@
On the other hand, by 4 in Appendix B, any surplus outcome $(\dot{P},\dot{C},\dot{S})$ satisfying $\frac{1}{2}\dot{P}+\dot{C}=\dot{S}$ can be achieved.
Part 2: Note the setting here is different from Part 1: in that case, we were constructing an environment without aggregate structure, while here we will give an example of an environment with recoverable structure where no intervention can robustly increase consumer surplus.
@@
We will study the distribution of this random variable for any fixed $\bm{\sigma}$, under the prior. We claim that $\dot{C}_{\geq 2}$ has the same distribution as $-\dot{C}_{\geq 2}$. This is because in the above expression, only $\bm{q}^{0}$ is random, and $\bm{u}^{\ell}\cdot\bm{q}^{0}$ are independently equal to plus or minus the same constant by definition. Thus, for any $\epsilon>0$,
@@
On the other hand, again by 4 in Appendix B, any surplus outcome $(\dot{P},\dot{C},\dot{S})$ satisfying $\frac{1}{2}\dot{P}+\dot{C}=\dot{S}$ can be achieved.
@@
For simplicity, this proof has worked with a case where all uncertainty comes from $\bm{q}^{0}$, and the projections of $\bm{q}^{0}$ onto various low eigenvectors are small. A stronger negative result can be obtained by elaborating the construction a bit—making the $\lambda_{\ell}$ different from one another, so that the eigenspaces orthogonal to $\bm{1}$ are not deterministic. Then, by showing that the eigenvectors are often impossible to recover with any precision, we can obtain a similar conclusion even if all the projections $\bm{u}^{\ell}\cdot\bm{q}^{0}$ are bounded below.
@@ Appendix B Complete information
First, consider the case in which products are independent. In this case, the effect of a subsidy to firm $i$ and a subsidy to firm $j$ to surpluses are separable. So it is sufficient to prove the result for an intervention which only subsidizes one firm: take $\bm{\sigma}$ with $\sigma_{j}=0$ for all $j\neq 1$ and $\sigma_{1}>0$, so that $\dot{S}=\sigma_{1}q_{1}^{0}$. Note the basis diagonalizing $\bm{D}$ is the standard basis, and using the formulas in 3 with $\lambda_{\ell}=-1$ for all $\ell$, we conclude $\dot{C}=\sigma q_{1}^{0}/2$ and $\dot{P}=\sigma q_{1}^{0}$, as claimed.
@@ Appendix C A noise structure under which 4 holds
For simplicity, we assume that all households share a single representative utility function for goods and that the number of households exceeds $n^{2}$, where $n$ is the number of firms. Moreover, we assume that, for all $i$ and $j$, the entries $\underline{D}_{ij}$ are uniformly bounded, i.e., $|\underline{D}_{ij}|\leq\underline{D}_{\operatornamewithlimits{max}}<\infty$ for some positive constant $\underline{D}_{\operatornamewithlimits{max}}$. Also assume the true diagonal entries $\underline{D}_{ii}$ are bounded away from 00: $0<\underline{D}_{\operatornamewithlimits{min}}\leq|\underline{D}_{ii}|$ for all $i$. The constants $\underline{D}_{\operatornamewithlimits{min}}$ and $\underline{D}_{\operatornamewithlimits{max}}$ do not depend on $i,j$, or $n$.
@@
We assume that the $F_{ij}$ are mean-zero and independent across $(i,j)$, that their support is uniformly bounded, and that the matrix $\bm{F}$ is symmetric. [^38] Since $\bm{D}$ is a symmetric matrix, we view the parameters as being the numbers on or above the diagonal. Let $\overline{V}$ be an upper bound on $\text{Var}[F_{ij}]$ that does not depend on $i,j$, or $n$.
@@
By standard results on the expected spectral norm of symmetric random matrices with independent entries (Bai and Yin, 1988), we have that $\|\bm{H}\|]=O_{p}(\sqrt{n})$. Under our assumption that $\bm{F}$ has uniformly bounded support, this can be extended (Dallaporta, 2012, Theorem 2) to the statement that $\mathbb{E}[\|\bm{H}\|]=O(\sqrt{n})$.